OpenStudy (anonymous):

file:///C:/Users/Parent/Downloads/ALGTests_L2_S1_U7_Q12v1Stem.gif

OpenStudy (anonymous):

A. y – 3 = –½(x + 1) B. y = ½x + 5 C. y – 3 = ½(x + 1) D. y + 3 = –½(x – 1)

OpenStudy (anonymous):

the link doesnt work